OpenStudy (anonymous):

You flip a coin 5 times. What is the probability that the results are all heads or all tails? Write your answer as a fraction

OpenStudy (the_fizicx99):

Well there's 2 sides to a coin so if you flip is 5 times, the probability is 1/2^5 or 1/32

OpenStudy (ybarrap):

So 1/2^5 is for all Heads add to this the probability of all tails, that is also 1/2^5. So probability of one OR the other, which are completely independent events, is 2 * 1/2^5 = 1/2^4 = 1/16.